OpenStudy (anonymous):

(x-4)-5(x-4)^1/2=6 solve for x.

OpenStudy (anonymous):

LET (x-1)^1/2 =y

OpenStudy (anonymous):

Then solve for y

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

not really...

OpenStudy (anonymous):

let \[x-4=m\] the equation becomes \[\left( m \right) - 5m^\frac{ 1 }{ 2 } = 6\] on squaring \[m^2 - 5m = 6\]

OpenStudy (anonymous):

therefore \[m=2 , m=2\]

OpenStudy (anonymous):

\[m=2 , m=3\] sry now as \[x-4 = m\] we get \[x-4=2 & x-4=3\]

OpenStudy (anonymous):

therefore x=6 and x =7

OpenStudy (anonymous):

i understand.. thank you!

OpenStudy (anonymous):

wheres my medal :D

OpenStudy (anonymous):

gotcha!!

OpenStudy (anonymous):

m^2 - 5m = 36 m^2 - 5m - 36 = 0 (m - 9)(m + 4) = 0 m = 9 and m = -4

OpenStudy (anonymous):

one thing that was missed was the 6. it wasnt squared.

OpenStudy (anonymous):

ya figured it out @franciscanmonk @bevandre sorry

OpenStudy (anonymous):

its all good.

OpenStudy (anonymous):

thanks i just needed someone to get me started...but i get it now:)

OpenStudy (anonymous):

If the problem expression is:\[(x - 4) - 5*\left(\sqrt{x-4}\right) = 6 \]then x = 40